OpenStudy (anonymous):

What are the factors of x2 - 16 ?

OpenStudy (anonymous):

Answer: a. (x+4)(x-4) b. (x-4)(x-4) c. (x+4)(x+4) d. x(x -4) e. (x^2-16)(x^2+16)

OpenStudy (raden):

hint : a^2 - b^2 = (a+b)(a-b)

OpenStudy (anonymous):

so is it e

OpenStudy (raden):

still incorrect

OpenStudy (raden):

u can rewrite x^2 - 16 = x^2 - 4^2

OpenStudy (anonymous):

hes right~!

OpenStudy (nubeer):

there is another way.. try to multiply you have in options.. which ever options turns out to be same as your question that will be your answer.. but lol not so reliable method

OpenStudy (raden):

but work too :)

OpenStudy (anonymous):

srry i dont get it guys will you just tell me the answer to it just this one time cause i dont get it

OpenStudy (nubeer):

ok this one time.. its A.. but you should still figure out the way how its done.. as everyone haev told and explained you above..

OpenStudy (anonymous):

i know but its hard im sorry guys i really am im on my own and only 17 and no one to help me

OpenStudy (nubeer):

hmm i am 19 and my own. just try to more of the practise. and you can always learn from here.. everyone is willling to help here who wants to learn..